Problem: From the information below, write an equation and then solve the equation.
A rectancle is 8cm longer than it is wide. The perimeter of the rectangle is 72 cm. Find the dimension of the rectangle.

Answer by checkley75(3666)

(w+8)=length
2w+2l=72 thus
2w+2(w+8)=72
2w=2w+16=72
4w=72-16
4w=56
w=56/4
w=14 answer forthe width
l=14+8
l=22 answer for the length.
proof
2*14=2*22=72
28+44=72
72=72